OpenStudy (anonymous):

The top of a tuna can has a circumference of 6 inches. Which lid will fit the top of the can? a. 1inch diameter b. 1 inch radius c. 2 inches diameter d. 2 inches radius

OpenStudy (anonymous):

C=pi*d 6=3.14d 6/3.14 = d 1.9 or 2 This is what I got. Please confirm.

OpenStudy (anonymous):

circumference= diametr * pi or 2* radius* pi

OpenStudy (paxpolaris):

that'a correct ... c. 2 inches diameter

OpenStudy (zepp):

\(C=2\pi r=6 \text{in}\) Therefore, \(2 \pi r=6 in\\\large r=\frac{6}{2 \pi}\approx0.95\)

OpenStudy (zepp):

So yes, you answer is right :D

OpenStudy (paxpolaris):

yes .. b and c are the same

OpenStudy (anonymous):

If its .95 that would be 1 inch, correct?

OpenStudy (zepp):

Yes, and that's the radius.

OpenStudy (anonymous):

I'm confused. I chose 2 inch diameter. Was that not correct?

OpenStudy (zepp):

A) and B) are both correct.

OpenStudy (anonymous):

it has to be one or the other. Which is the BEST one?

OpenStudy (paxpolaris):

B and C are equivalent ... neither is better than the other

OpenStudy (zepp):

B and C, sorry

OpenStudy (zepp):

I would say if there's only one answer, this question fails, lol :P